# Exploring Weight Balancing on Long-Tailed Recognition Problem

### v5

@@ A.3.2 Proof of Theorem 2
For any $k\in\mathcal{Y}$, if there exists $\mathbf{w}_{k}^{*}$ s.t. $\left.\frac{\partial F_{\mathrm{WB}}}{\partial\mathbf{w}_{k}}\right|_{\mathbf{w}_{k}=\mathbf{w}_{k}^{*}}=\mathbf{0}$ and $\|\mathbf{w}_{k}^{*}\|_{2}{\color[rgb]{1,0,0}=}O\left(\frac{1}{\lambda\rho C}\right)$, we have
@@ A.4 Settings
Our implementation is based on Alshammari et al. [2022] and Vigneswaran et al. [2021]. For comparison, many of the experimental settings also follow Alshammari et al. [2022].
@@ A.4.1 Datasests
We created validation datasets from the portions of the training datasets because CIFAR10 and CIFAR100 have only training and test data. As with Liu et al. [2019], only $20$ samples per class were taken from the training dataset to compose the validation dataset, and the training dataset was composed of the rest of the data. We set $N_{1}$ to $4980$ for CIFAR10 and $480$ for CIFAR100. In our experiments, we set imbalance factor $\rho$ to $100$. We call the class $k$ Many if the number of training samples satisfies $1000<N_{k}\leq 4980$ (resp. $100<N_{k}\leq 480$), Medium if the number of training samples fullfills $200\leq N_{k}\leq 1000$ (resp. $20\leq N_{k}\leq 100$), and Few otherwise for CIFAR10-LT (resp. CIFAR100-LT). For mini-ImageNet-LT and ImageNet-LT, the same applies as for CIFAR100-LT.
@@ A.4.2 Models
As for MLPs, one module block consists of three layers: a linear layer outputting $1024$ dimensional features, a BN layer, and a ReLU layer. The layers are stacked in sequence and the blocks composed of them are as well. As for ResBlocks, each block has the same structure as He et al. [2016], except that the linear layer outputs $1024$ dimensional features. These blocks are combined into a sequential, and at the bottom of it, we further concatenate an MLP block. In other words, the input first passes through the linear layer, BN, and ReLU before flowing into the residual blocks. In both cases, a classifier consisting of a linear layer and a softmax activation layer is on the top and does not count as one block. For example, in MLP3, the features pass through the three blocks and the classifier in sequence.
@@ A.4.3 Evaluation metrics
For ImageNet-LT, We searched hyperparameters using the validation dataset for the ones not published in Alshammari et al. [2022] and used the following values. We reduced the learning rate gradually from $0.05$ to 00. The number of epochs was $200$ for the first stage. We set $\lambda$ of WD to $0.00024$ in the first stage and $0.00003$ in the second stage. We set $\zeta$ of FR to $0.0001$. The other hyperparameters were the same as in CIFAR100-LT.
@@ A.4.4 Why we do not use DR loss
This section explains why we use ETF classifiers but not the dot-regression loss (DR) proposed in Yang et al. [2022]. Table 9 compares FDRs and accuracy when we use WD and an ETF Classifier for the linear layer with two options, using CE or DR for the loss. This table shows that CE is superior in the test FDR and average accuracy. Theorem 1 also claims that training by CE and WD prevents the cone effect. It does not apply to DR, which is a squared loss function. Thus, it is difficult to prove whether DR is equally effective in preventing the cone effect.
@@ A.5 Experiments
The right half of Figure 4 show the result of Sec.4.2 for mini-ImageNet-LT. We also examined the forget score [Toneva et al., 2019] of each method in Sec.4.2. Phenomena similar to the ones we showed in Sec.4.2 can be observed in the forgetting score of features trained with CB: Figure 3 shows that the forgetting score is higher on average with CB than with CE. In the absence of WD, this phenomenon is particularly evident for the Many classes, possibly because CB gives each image in Many classes less weight during training, which inhibits learning. These may be some of the reasons for the poor accuracy of CB compared with CE in training feature extractors reported by Kang et al. [2020].
@@
First, we experimentally show MaxNorm is not necessary for the second stage of WB. In the second training stage in WB, we observed that normalizing the norm of the weights to one before training (WD + WD & Norm & CB) instead of applying MaxNorm every epoch gives almost identical results as shown in Figure 5. This phenomenon is consistent with the fact that the regularization loses its effect beyond a certain epoch [Golatkar et al., 2019].
@@
We also experimented with the ResNeXt50 [Xie et al., 2017] as with Alshammari et al. [2022]. We used the same hyperparameters as the ResNet. Table 15 presents the results for CIFAR100-LT.
@@ A.5.6 Experiments on Tabular Data
Kadra et al. [2021] show that even an MLP with carefully tuned regularization outperforms state-of-the-art models in the classification of tabular data. Following them, we used a 9-layer MLP with $512$ dimensions per layer for the feature extractor. The model is trained by AdamW [Loshchilov and Hutter, 2018] for $400$ epochs in the first stage and $10$ epochs in the second stage. Thus, WD is not implemented in L2 regularization in this dataset, but is built into the optimizer. By a hyperparameter search with the validation data, we set the hyperparameters as follows. We used dropout [Srivastava et al., 2014] and set the hyperparameter to $0.15$. The initial learning rate was $0.001$ for the first stage and $0.0004$ for the second stage. We set $\lambda$ of WD to $0.15$ for the first stage and $0.0003$ for the second stage. We set $\zeta$ of FR to $0.001$.
@@ A.6 Broader impacts
Our research provides theoretical and experimental evidence for the effectiveness of an existing ad-hoc method. We also show that the original method can be simplified based on this theory. The theory is valid for general deep neural networks and does not concern any social issues such as privacy. On the contrary, it reduces the number of training stages to one while maintaining a higher level of accuracy, thus reducing the computational cost and the negative impact on the environment.

### v7

@@ A.3.2 Proof of Theorem 2
For any $k\in\mathcal{Y}$, if there exists $\mathbf{w}_{k}^{*}$ s.t. $\left.\frac{\partial F_{\mathrm{WB}}}{\partial\mathbf{w}_{k}}\right|_{\mathbf{w}_{k}=\mathbf{w}_{k}^{*}}=\mathbf{0}$ and $\|\mathbf{w}_{k}^{*}\|_{2}=O\left(\frac{1}{\lambda\rho C}\right)$, we have
@@ A.4 Settings
Our implementation is based on Alshammari et al. (2022) and Vigneswaran et al. (2021). For comparison, many of the experimental settings also follow Alshammari et al. (2022).
@@ A.4.1 Datasests
We created validation datasets from the portions of the training datasets because CIFAR10 and CIFAR100 have only training and test data. As with Liu et al. (2019), only $20$ samples per class were taken from the training dataset to compose the validation dataset, and the training dataset was composed of the rest of the data. We set $N_{1}$ to $4980$ for CIFAR10 and $480$ for CIFAR100. In our experiments, we set imbalance factor $\rho$ to $100$. We call the class $k$ Many if the number of training samples satisfies $1000<N_{k}\leq 4980$ (resp. $100<N_{k}\leq 480$), Medium if the number of training samples fullfills $200\leq N_{k}\leq 1000$ (resp. $20\leq N_{k}\leq 100$), and Few otherwise for CIFAR10-LT (resp. CIFAR100-LT). For mini-ImageNet-LT and ImageNet-LT, the same applies as for CIFAR100-LT.
@@ A.4.2 Models
As for MLPs, one module block consists of three layers: a linear layer outputting $1024$ dimensional features, a BN layer, and a ReLU layer. The layers are stacked in sequence and the blocks composed of them are as well. As for ResBlocks, each block has the same structure as He et al. (2016), except that the linear layer outputs $1024$ dimensional features. These blocks are combined into a sequential, and at the bottom of it, we further concatenate an MLP block. In other words, the input first passes through the linear layer, BN, and ReLU before flowing into the residual blocks. In both cases, a classifier consisting of a linear layer and a softmax activation layer is on the top and does not count as one block. For example, in MLP3, the features pass through the three blocks and the classifier in sequence.
@@ A.4.3 Evaluation metrics
For ImageNet-LT, We searched hyperparameters using the validation dataset for the ones not published in Alshammari et al. (2022) and used the following values. We reduced the learning rate gradually from $0.05$ to 00. The number of epochs was $200$ for the first stage. We set $\lambda$ of WD to $0.00024$ in the first stage and $0.00003$ in the second stage. We set $\zeta$ of FR to $0.0001$. The other hyperparameters were the same as in CIFAR100-LT.
@@ A.4.4 Why we do not use DR loss
This section explains why we use ETF classifiers but not the dot-regression loss (DR) proposed in Yang et al. (2022). Table 9 compares FDRs and accuracy when we use WD and an ETF Classifier for the linear layer with two options, using CE or DR for the loss. This table shows that CE is superior in the test FDR and average accuracy. Theorem 1 also claims that training by CE and WD prevents the cone effect. It does not apply to DR, which is a squared loss function. Thus, it is difficult to prove whether DR is equally effective in preventing the cone effect.
@@ A.5 Experiments
The right half of Figure 4 show the result of Sec.4.2 for mini-ImageNet-LT. We also examined the forget score (Toneva et al., 2019) of each method in Sec.4.2. Phenomena similar to the ones we showed in Sec.4.2 can be observed in the forgetting score of features trained with CB: Figure 3 shows that the forgetting score is higher on average with CB than with CE. In the absence of WD, this phenomenon is particularly evident for the Many classes, possibly because CB gives each image in Many classes less weight during training, which inhibits learning. These may be some of the reasons for the poor accuracy of CB compared with CE in training feature extractors reported by Kang et al. (2020).
@@
First, we experimentally show MaxNorm is not necessary for the second stage of WB. In the second training stage in WB, we observed that normalizing the norm of the weights to one before training (WD + WD & Norm & CB) instead of applying MaxNorm every epoch gives almost identical results as shown in Figure 5. This phenomenon is consistent with the fact that the regularization loses its effect beyond a certain epoch (Golatkar et al., 2019).
@@
We also experimented with the ResNeXt50 (Xie et al., 2017) as with Alshammari et al. (2022). We used the same hyperparameters as the ResNet. Table 15 presents the results for CIFAR100-LT.
@@ A.5.6 Experiments on Tabular Data
Kadra et al. (2021) show that even an MLP with carefully tuned regularization outperforms state-of-the-art models in the classification of tabular data. Following them, we used a 9-layer MLP with $512$ dimensions per layer for the feature extractor. The model is trained by AdamW (Loshchilov & Hutter, 2018) for $400$ epochs in the first stage and $10$ epochs in the second stage. Thus, WD is not implemented in L2 regularization in this dataset, but is built into the optimizer. By a hyperparameter search with the validation data, we set the hyperparameters as follows. We used dropout (Srivastava et al., 2014) and set the hyperparameter to $0.15$. The initial learning rate was $0.001$ for the first stage and $0.0004$ for the second stage. We set $\lambda$ of WD to $0.15$ for the first stage and $0.0003$ for the second stage. We set $\zeta$ of FR to $0.001$.
@@ A.6 Broader impacts
Our research provides theoretical and experimental evidence for the effectiveness of an existing ad-hoc method. We also show that the original method can be simplified based on this theory. The theory is valid for general DNNs and does not concern any social issues such as privacy. On the contrary, it reduces the number of training stages to one while maintaining a higher level of accuracy, thus reducing the computational cost and the negative impact on the environment.